The temperature coefficient for copper is 0.0039 (◦C)^−1 . If the wire is heated to 36◦C while the 6 V potential difference is maintained, what is the resulting current in the wire? Answer in units of A.
Answer #1

The typical values of copper at 20 C with initial resistance 100 ohms are:

R=R_{0}(1+\alpha \Delta T)=100\Omega (1+0.0039\cdot (36C^{\circ}-20C^{\circ}))=106.24\Omega

FRom ohms law:

V=IR

I=\frac{V}{R}=\frac{6V}{106.24}=0,056\, A